OpenStudy (anonymous):

What two numbers give me a product of 75 and a sum of -28

OpenStudy (anonymous):

xy=75 and x+y=-28 Rearrange the second equation to get x=-y-28. Substitute this into the first equation to get -y(y+28)=75. Expand and take over to RHS: y^2 +28y+75=0. Solve using quadratic equation: y=-3 and -25 Substitute this back into the second equation: When y=-3, x=-25 When y=-25, x=3

OpenStudy (anonymous):

negative 25 and negative 3

OpenStudy (anonymous):

Not sure how to edit my post, but I made a typo at the end. (It should be y=-25 and x=-3 again).